# Approximating Two-Layer ReLU Networks for Hidden State Analysis in Differential Privacy

### v2

@@ Abstract
The hidden state threat model of differential privacy (DP) assumes that the adversary has access only to the final trained machine learning (ML) model, without seeing intermediate states during training. Current privacy analyses under this model, however, are limited to convex optimization problems, reducing their applicability to multi-layer neural networks, which are essential in modern deep learning applications. Additionally, the most successful applications of the hidden state privacy analyses in classification tasks have been for logistic regression models. We demonstrate that it is possible to privately train convex problems with privacy-utility trade-offs comparable to those of one hidden-layer ReLU networks trained with DP stochastic gradient descent (DP-SGD). We achieve this through a stochastic approximation of a dual formulation of the ReLU minimization problem which results in a strongly convex problem. This enables the use of existing hidden state privacy analyses, providing accurate privacy bounds also for the noisy cyclic mini-batch gradient descent (NoisyCGD) method with fixed disjoint mini-batches. Our experiments on benchmark classification tasks show that NoisyCGD can achieve privacy-utility trade-offs comparable to DP-SGD applied to one-hidden-layer ReLU networks. Additionally, we provide theoretical utility bounds that highlight the speed-ups gained through the convex approximation.
@@ 1 Introduction
In differentially private (DP) machine learning (ML), the DP-SGD algorithm (see e.g., Abadi et al.,, 2016) has become a standard tool obtain ML models with strong privacy guarantees of the individuals. The $(\varepsilon,\delta)$-guarantees for DP-SGD are obtained by clipping gradients by adding normally distributed noise scaled with the clipping constant to the randomly sampled mini-batch of gradients, and by using composition DP analysis (see, e.g., Zhu et al.,, 2022).
@@
One weak point of the composition analysis of DP-SGD is that it is assumes that the adversary has access to all the intermediate results of the training iteration. This assumption is often unnecessarily strict as in many practical scenarios only the final model is needs to be revealed. Another weakness is that it requires either full batch training or random subsampling, and, e.g., accurate privacy analyses for of many practically relevant algorithms are not available for non-convex problems such as for training multi-layer neural networks. One example of such algorithms is the noisy cyclic GD (NoisyCGD) with disjoint mini-batches. Having high privacy-utility ML models trained with NoisyCGD would give an alternative for DP-SGD that is often difficult to implement in practical settings (Chua et al.,, 2024).
@@
The so called hidden state threat model of DP considers releasing only the final model of the training iteration, and existing $(\varepsilon,\delta)$-DP analyses in the literature are only applicable for convex problems such as the logistic regression which is the highest performing model considered in the literature (see, e.g., Chourasia et al.,, 2021; Bok et al.,, 2024). When training models with DP-SGD, however, one quickly finds that the model performance of commonly used convex models is inferior compared to multi-layer neural networks. A natural question then arises whether convex approximations of minimization problems for multi-layer neural networks can be made while preserving model performance. In response, this work explores such an approximation for the two-layer ReLU minimization problem. To achieve this, we build on the findings of Pilanci and Ergen, (2020), which demonstrate the existence of a convex dual formulation for the two-layer ReLU minimization problem when the hidden layer is sufficiently wide.
@@
The privacy amplification by iteration analysis for convex private optimization, introduced by Feldman et al., (2018), provides privacy guarantees in the hidden state threat model. However, this and many subsequent analyses (Sordello et al.,, 2021; Asoodeh et al.,, 2020; Chourasia et al.,, 2021; Altschuler and Talwar,, 2022) remain challenging to apply in practice, as it typically requires a large number of training iterations for obtaining tighter DP guarantees than those of DP-SGD. Chourasia et al., (2021) improved this analysis using Rényi DP for full-batch DP-GD training, while Ye and Shokri, (2022) offered a similar analysis for shuffled mini-batch DP-SGD. Recently, Bok et al., (2024) provided an $f$-DP analysis for a class of algorithms, which we also leverage to analyze NoisyCGD.
Also from a theoretical standpoint, convex models are advantageous over non-convex ones in private optimization. State-of-the-art empirical risk minimization (ERM) bounds for private convex optimization are of the order $O(\tfrac{1}{\sqrt{n}}+\tfrac{\sqrt{d}}{\varepsilon n})$, where $n$ is the number of training data entries, $d$ the dimension of the parameter space and $\varepsilon$ the DP parameter (Bassily et al.,, 2019). In contrast, the bounds for non-convex optimization, which focus on finding stability points, are much worse, such as $O(\tfrac{1}{n^{1/3}}+\tfrac{d^{1/5}}{(\varepsilon n)^{2/5}})$ (Bassily et al.,, 2021) and $O\left(\tfrac{d^{1/3}}{(\varepsilon n)^{2/3}}\right)$ (Arora et al.,, 2023; Lowy et al.,, 2024).
@@
  • By integrating two seemingly unrelated approaches, convex reformulation of ReLU networks and privacy amplification by iteration DP analysis, we show that it is possible to obtain similar privacy-utility trade-offs in the hidden state threat model of DP as by applying DP-SGD to two-layer ReLU networks and using composition results.
  • We carry out a number of approximations for the convex reformulation to facilitate DP analysis and show that the resulting strongly convex model has the required properties for hidden state analysis.
  • We give the first high privacy-utility trade-off results for benchmark classification tasks using a hidden state DP analysis. In particular, we give the first empirical high privacy-utility trade-off results for NoisyCGD with disjoint mini-batches under the hidden state threat model which makes it more suitable for practical applications of DP ML. The experiments also account for the privacy cost of hyperparameter tuning, and we demonstrate how to conduct it effectively for NoisyCGD.
  • We carry out a theoretical utility analysis of DP-SGD applied to the convex approximation within the random data model.
@@ 2 Preliminaries
We denote a dataset containing $n$ data points as $D=(z_{1},\ldots,z_{n}).$ We say $D$ and $D^{\prime}$ are neighboring datasets if they differ in exactly one element (denoted as $D\sim D^{\prime}$). We say that a mechanism $\mathcal{M}\,:\,\mathcal{X}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable (Dwork et al.,, 2006).
@@
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. Mechanism $\mathcal{M}\,:\,\mathcal{X}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if for every pair of neighboring datasets $D,D^{\prime}$ and for every measurable set $E\subset\mathcal{O}$,
@@
Hockey-stick Divergence and Numerical Privacy Accounting.
The DP guarantees can be alternatively described using the hockey-stick divergence which is defined as follows. For $\alpha>0$ the hockey-stick divergence $H_{\alpha}$ from a distribution $P$ to a distribution $Q$ is defined as
$H_{\alpha}(P||Q)=\int\left[P(t)-\alpha\cdot Q(t)\right]_{+}\,\hskip 1.0pt{\rm d}\hskip 0.5ptt,$ (2.1)
where for $t\in\mathbb{R}$, $[t]_{+}=\max\{0,t\}$. The $(\varepsilon,\delta)$-DP guarantee as defined in Def. 2.1 can be characterized using the hockey-stick divergence as follows.
Lemma 2.2 (Zhu et al., 2022).
For a given $\varepsilon\geq 0$, tight $\delta(\varepsilon)$ is given by the expression
$\delta(\varepsilon)=\max_{D\sim D^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(D)||\mathcal{M}(D^{\prime})).$
Thus, if we can bound the divergence $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(D)||\mathcal{M}(D^{\prime}))$ accurately, we also obtain accurate $\delta(\varepsilon)$-bounds. We also refer to $\delta_{\mathcal{M}}(\varepsilon):=\max_{D\sim D^{\prime}}H_{e^{\epsilon}}(\mathcal{M}(D)||\mathcal{M}(D^{\prime}))$ as the privacy profile of mechanism $\mathcal{M}$. For bounding the hockey-stick divergence of compositions accurately, we need to so-called dominating pairs of distributions.
@@
If the equality holds for all $\alpha$ for some $D,D^{\prime}$, then $(P,Q)$ is a tightly dominating pair of distributions. We get upper bounds for DP-SGD compositions using the dominating pairs of distributions using the following composition result.
@@ 2.1 DP-SGD with Poisson Subsampling
DP-SGD iteration with Poisson subsampling is given by
$\theta_{j+1}=\theta_{j}-\eta_{j}\cdot\Big{(}\frac{1}{b}\sum\nolimits_{x\in B_{j}}\text{clip}(\nabla\mathcal{L}(x,\theta_{j}),C)+Z_{j}\Big{)},$ (2.2)
where $C>0$ denotes the clipping constant, $\text{clip}(\cdot,C)$ the clipping function that clips gradients to have 2-norm at most $C$, $\mathcal{L}$ the loss function, $\theta$ the model parameters, $\eta_{j}$ the learning rate at iteration $j$, $B_{j}$ the mini-batch at iteration $j$ that is sampled with Poisson subsampling with the subsampling ratio $b/n$, $b$ the expected size of each mini-batch and $Z_{j}\sim\mathcal{N}(0,\tfrac{C^{2}\sigma^{2}}{b^{2}}I_{d})$ the noise vector.
We want to experimentally compare DP-SGD to the noisy cyclic mini-batch gradient descent using the privacy amplification by iteration analysis Bok et al., (2024). To this end, we consider the substitute neighborhood relation of datasets. To this end, we use to following results by Lebeda et al., (2024).
@@
Suppose a pair of distributions $(P,Q)$ is a dominating pair of distributions for a mechanism $\mathcal{M}$ and denote the Poisson subsampled mechanism ~$\widetilde{\mathcal{M}}:=\mathcal{M}\circ S^{q}_{Poisson}$, where $S^{q}_{Poisson}$ denotes the Poisson subsampling with subsampling ratio $q$. Then, for all neighbouring datasets (under the $\sim$-neighbouring relation) $X$ and $Y$,
$H_{\alpha}\big{(}\widetilde{\mathcal{M}}(X)||\widetilde{\mathcal{M}}(Y)\big{)}$ (2.3)
$\leq H_{\alpha}\big{(}(1-q)\cdot\mathcal{N}(0,\sigma^{2})+q\cdot\mathcal{N}(1,\sigma^{2})||(1-q)\cdot\mathcal{N}(0,\sigma^{2})+q\cdot\mathcal{N}(-1,\sigma^{2})\big{)}$
for all $\alpha\geq 0$.
Furthermore, using the composition result of Lemma 2.4 and numerical accountants, we obtain $(\varepsilon,\delta)$-bounds for compositions of DP-SGD with Poisson subsampling the substitute neighborhood relation of datasets. Alternatively, we could use RDP bounds given by Wang et al., (2019), however, as also illustrated by the Appendix Figure 6, our numerical approach generally leads to tighter bounds.
2.2 Guarantees for the Final Model and for Noisy Cyclic Mini-Batch GD
We next consider privacy amplification by iteration (Feldman et al.,, 2018) type of analysis that gives DP guarantees for the final model of the training iteration. We use the recent results by Bok et al., (2024) that are applicable to the noisy cyclic mini-batch gradient descent (NoisyCGD) for which one epoch of training is described by the iteration
$\theta_{j+1}=\theta_{j}-\eta\left(\frac{1}{b}\sum\nolimits_{x\in B_{j}}\nabla_{\theta}f(\theta_{j},x)+Z_{j}\right)$ (2.4)
where $Z_{j}\sim\mathcal{N}(0,\sigma^{2}I_{d})$ and the data $D$, $\left|D\right|=n$, is divided into disjoint batches $B_{1},\ldots,B_{k}$, each of size $b$. The analysis by Bok et al., (2024) considers the substitute neighborhood relation of datasets and central for the DP analysis is the gradient sensitivity.
@@ 2.2 Guarantees for the Final Model and for Noisy Cyclic Mini-Batch GD
We say that a family of loss functions $\mathcal{F}$ has a gradient sensitivity $L$ if
$\sup_{f,g\in\mathcal{F}}\left\lVert\nabla f-\nabla g\right\rVert\leq L.$
As an example relevant to our analysis, for a family of loss functions of the form $h_{i}+r$, where $h_{i}$’s are $L$-Lipschitz loss functions and $r$ is a regularization function, the sensitivity equals $2L$. We will use the following result for analysing the $(\varepsilon,\delta)$-DP guarantees of NoisyCGD. First, recall that a function $f$ is $\beta$-smooth if $\nabla f$ is $\beta$-Lipschitz, and it is $\lambda$-strongly convex if the function $g(x)=f(x)-\frac{\lambda}{2}\left\lVert x\right\rVert_{2}^{2}$ is convex.
@@
Consider $\lambda$-strongly convex, $\beta$-smooth loss functions with gradient sensitivity $L$. Then, for any $\eta\in(0,2/M)$, NoisyCGD is $\mu$-GDP for
@@
We could alternatively use the RDP analysis by Ye and Shokri, (2022), however, as also illustrated by the experiments of Bok et al., (2024), the bounds given by Thm. 2.8 lead to slightly lower $(\varepsilon,\delta)$-DP bounds for NoisyCGD.
In order to benefit from the privacy analysis of Thm. 2.8 for NoisyCGD, we add an $L_{2}$-regularization term with a coefficient $\tfrac{\lambda}{2}$ which makes the loss function $\lambda$-strongly convex. Finding suitable values for the learning rate $\eta$ and regularization parameter $\lambda$ is complicated by the following aspects. The larger the regularization parameter $\lambda$ and the learning rate $\eta$ are, the faster the model ’forgets’ the past updates and the faster the $\varepsilon$-values converge. This is reflected in the GDP bound of Thm.  2.8 in the constant $c$ which generally equals $\left|1-\eta\lambda\right|$. Thus, in order to benefit from the bound of Thm.  2.8, the product $\eta\lambda$ should not be too small. On the other hand, when $\eta\lambda$ is too large, the ’forgetting’ starts to affect the model performance. We experimentally observe that the plateauing of the model accuracy and privacy guarantees happens approximately at the same time.
Figure 1 illustrates privacy guarantees of NoisyCGD for a range of values for the product $\eta\lambda$ where the $(\varepsilon,\delta)$-DP guarantees given by Thm.  2.8 become smaller than those given by the Poisson subsampled DP-SGD with an equal batch size $b=1000$ when $\sigma=15.0$ when training for 400 epochs. For a given value of the learning $\eta$, we can always adjust the value of $\lambda$ to have desirable $(\varepsilon,\delta)$-DP guarantees. To put the values of Fig. 1 into perspective, in experiments we observe that $\eta\lambda=2\cdot 10^{-4}$ is experimentally found to already affect the model performance considerably whereas $\eta\lambda=1\cdot 10^{-4}$ affects only weakly.
Figure 1: Values of the product of the learning rate $\eta$ and the $L_{2}$-regularization constant $\lambda$ that lead to tighter privacy bounds for the final model using Thm. 2.8 than for the whole sequence of updates using the DP-SGD analysis. Here the $\varepsilon$-values are shown as a function of number of training epochs, when total number of training samples $N=6\cdot 10^{4}$, the batch size $b=1000$, $\sigma=15.0$ and $\delta=10^{-5}$.

### v3

# Convex Approximation of Two-Layer ReLU Networks for Hidden State Differential Privacy

@@ Abstract
The hidden state threat model of differential privacy (DP) assumes that the adversary has access only to the final trained machine learning (ML) model, without seeing intermediate states during training. However, the current privacy analyses under this model are restricted to convex optimization problems, reducing their applicability to multi-layer neural networks, which are essential in modern deep learning applications. Notably, the most successful applications of the hidden state privacy analyses in classification tasks have only been for logistic regression models. We demonstrate that it is possible to privately train convex problems with privacy-utility trade-offs comparable to those of 2-layer ReLU networks trained with DP stochastic gradient descent (DP-SGD). This is achieved through a stochastic approximation of a dual formulation of the ReLU minimization problem, resulting in a strongly convex problem. This enables the use of existing hidden state privacy analyses and provides accurate privacy bounds also for the noisy cyclic mini-batch gradient descent (NoisyCGD) method with fixed disjoint mini-batches. Empirical results on benchmark classification tasks demonstrate that NoisyCGD can achieve privacy-utility trade-offs on par with DP-SGD applied to 2-layer ReLU networks.
@@ 1 Introduction
In differentially private (DP) machine learning (ML), the DP-SGD algorithm (see e.g., Abadi et al., , 2016) has become a ubiquitous tool to obtain ML models with strong privacy guarantees. The guarantees for DP-SGD are obtained by clipping gradients and by adding normally distributed noise to the randomly sampled mini-batch of gradients. The parameters for the DP guarantee are then found by a composition analysis of the iterative algorithm (see, e.g., Zhu et al., , 2022).
@@
One weak point of the composition analysis of DP-SGD is the assumption that the adversary has access to all the intermediate results of the training iteration. This assumption is unnecessarily strict as in many practical scenarios only the final model needs to be revealed (Andrew et al., , 2024). Another weakness is that DP-SGD requires either full batch training or random subsampling. Both are computationally not favourable, as full batch usually does not fit in memory and random subsampling implies an unbalanced use of compute resources and increased implementation complexity (Chua et al., 2024a, ). In contrast, for noisy cyclic gradient descent (NoisyGD) with disjoint mini-batches, the compute is balanced between mini-batches and can be implemented efficiently. Having high privacy-utility ML models trained with NoisyCGD would give an alternative for DP-SGD that is more practical and efficient to implement (Chua et al., 2024a, ).
@@
The so-called hidden state threat model of DP considers releasing only the final model of the training iteration. Existing $(\varepsilon,\delta)$DP analyses in the literature are only applicable for convex problems such as logistic regression which is the highest performing model considered in the literature (see, e.g., Chourasia et al., , 2021; Bok et al., , 2024). When training models with DP-SGD, however, one quickly finds that the model performance of commonly used convex models is inferior compared to neural networks. A natural question arises whether convex approximations of minimization problems for multi-layer neural networks can be made while preserving model performance under privacy. This work explores such an approximation for the two-layer ReLU neural network. We build on the findings of Pilanci and Ergen, (2020), which demonstrate the existence of a convex dual formulation for the two-layer ReLU minimization problem when the hidden layer has sufficient width.
@@
The privacy amplification by iteration analysis for convex private optimization, introduced by Feldman et al., (2018), provides privacy guarantees in the hidden state threat model. However, these analyses (Sordello et al., , 2021; Asoodeh et al., , 2020; Chourasia et al., , 2021; Altschuler and Talwar, , 2022) remain challenging to apply in practice, as they typically require a large number of training iterations to obtain tighter DP guarantees than those of DP-SGD. Chourasia et al., (2021) improve this analysis using Rényi DP for full-batch training with DP guarantees, while Ye and Shokri, (2022) offer a similar analysis for shuffled mini-batch DP-SGD. Recently, Bok et al., (2024) provided an $f$-DP analysis for a class of algorithms, which we will leverage to analyze NoisyCGD.
@@
  • By integrating two seemingly unrelated approaches, a convex reformulation of ReLU networks and the privacy amplification by iteration DP analysis, we show that it is possible to obtain similar privacy-utility trade-offs in the hidden state threat model of DP as by applying DP-SGD to two-layer ReLU networks and using well-known composition results.
  • We provide approximations for the convex reformulation to facilitate DP analysis and show that the resulting strongly convex model has the required properties for hidden state analysis.
  • We give the first high-privacy-utility trade-off results for image classification tasks using a hidden state DP analysis. In particular, we present results for NoisyCGD with disjoint mini-batches. This allows for more practical applications of DP ML.
  • We run a theoretical utility analysis of gradient descent with DP guarantees when applied to the convex approximation, in the context of the random data model.
1.1 Further Related Literature
Differentially private machine learning has been extensively studied in recent years. The most widely used approach is DP-SGD (Abadi et al., , 2016), which clips and adds noise to gradients during training. Several works have analyzed and improved DP-SGD through techniques such as adaptive clipping (Andrew et al., , 2021), better composition analyses (Dong et al., , 2022; Koskela et al., , 2020; Zhu et al., , 2022; Gopi et al., , 2021), or privacy amplification by iteration (Feldman et al., , 2018). For convex problems, alternative approaches include sufficient statistics perturbation (Wang, , 2018; Amin et al., , 2023), objective perturbation (Chaudhuri et al., , 2011), and output perturbation (Wu et al., , 2017). Most prior work focuses on the standard DP definition rather than hidden state DP. Another alternative to DP-SGD is the Differentially Private Follow-the-Regularized-Leader (Kairouz et al., , 2021), but it comes with significant additional memory and compute cost due to sampling of correlated noise from a matrix mechanism, and potentially significant communication overhead in the federated learning setting. A recent line of work considers DP guarantees for DP-SGD with disjoint batches. Chua et al., (2025) andChoquette-Choo et al., (2025) use Monte Carlo methods to estimate privacy guarantees under the so-called balls-and-bins sampling, where Poisson sampling assigns data points to disjoint batches. However, these simulations are computationally intensive and do not lead to provable upper bounds on DP parameters. Feldman and Shenfeld, (2025) provide provable bounds for a generalization of this scheme, though their method also results in randomly sized batches. Thus, existing analyses do not accommodate fixed-size batches and in particular unshuffled data, further motivating our approach, which yields high-utility convex models for which NoisyCGD can be analyzed accurately.
Our work builds on recent advances in convex approximations of neural networks. The connection between two-layer ReLU networks and convex optimization was established by Bengio et al., (2005); Bach, (2017) and further developed by Pilanci and Ergen, (2020). Several works propose methods to train neural networks through convex optimization (Ergen and Pilanci, , 2020; Ergen et al., , 2023). However, the privacy implications of these convex formulations have not been thoroughly explored before our work. The closest related work is (Kim and Pilanci, , 2024), which analyzes connections between stochastic dual forms and ReLU networks but does not consider privacy.
@@ 2 Preliminaries
We denote a dataset containing $n$ data points as $D=(z_{1},\ldots,z_{n}).$ We say $D$ and $D^{\prime}$ are neighboring datasets if they differ in exactly one element (denoted as $D\sim D^{\prime}$). A mechanism $\mathcal{M}\,:\,\mathcal{X}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable (Dwork et al., , 2006).
@@
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. Mechanism $\mathcal{M}\,:\,\mathcal{X}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if for every pair of neighboring datasets $D\sim D^{\prime}$ and for every measurable set $E\subset\mathcal{O}$,
@@
The DP guarantees can alternatively be described using the hockey-stick divergence. For $\alpha>0$ the hockey-stick divergence $H_{\alpha}$ from a distribution $P$ to a distribution $Q$ is defined as $H_{\alpha}(P||Q)=\int\max\{P(t)-\alpha\cdot Q(t),0\}\,\hskip 1.0pt{\rm d}\hskip 0.5ptt.$ The $(\varepsilon,\delta)$-DP guarantee, in Def. 2.1, can be characterized using the hockey-stick divergence: if we can bound the divergence $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(D)||\mathcal{M}(D^{\prime}))$ accurately, we also obtain accurate $\delta(\varepsilon)$-bounds. We also refer to $\delta_{\mathcal{M}}(\varepsilon):=\max_{D\sim D^{\prime}}H_{e^{\epsilon}}(\mathcal{M}(D)||\mathcal{M}(D^{\prime}))$ as the privacy profile of mechanism $\mathcal{M}$. To accurately bound the hockey-stick divergence of compositions, we need to so-called dominating pairs of distributions.
@@
If the equality holds for all $\alpha$ for some $D\sim D^{\prime}$, then $(P,Q)$ is a tightly dominating pair of distributions. We get upper bounds for DP-SGD compositions using the dominating pairs of distributions using the following composition result.
@@ 2.1 DP-SGD with Poisson Subsampling
One iteration of DP-SGD with Poisson subsampling is given by
$\theta_{j+1}=\theta_{j}-\eta_{j}\cdot\left(\frac{1}{b}\sum\nolimits_{x\in B_{j}}\text{clip}(\nabla\mathcal{L}(x,\theta_{j}),C)+Z_{j}\right),$
where $C>0$ denotes the clipping constant, $\text{clip}(\cdot,C)$ the clipping function that clips gradients to have the 2-norm at most $C$, $\mathcal{L}$ the loss function, $\theta$ the model parameters, $\eta_{j}$ the learning rate at iteration $j$, $B_{j}$ the mini-batch at iteration $j$ sampled with Poisson subsampling with the subsampling ratio $b/n$, $b$ the expected size of each mini-batch and $Z_{j}\sim\mathcal{N}(0,\tfrac{C^{2}\sigma^{2}}{b^{2}}I_{d})$ the noise vector.
We adopt the substitute neighborhood relation and apply the result of Lebeda et al., (2024):
@@
Suppose a pair of distributions $(P,Q)$ is a dominating pair of distributions for a mechanism $\mathcal{M}$ and denote the Poisson subsampled mechanism ~$\widetilde{\mathcal{M}}:=\mathcal{M}\circ S^{q}_{Poisson}$, where $S^{q}_{Poisson}$ denotes the Poisson subsampling with subsampling ratio $q$. Then, under the $\sim$-neighbouring relation, the pair of distributions $(P,Q)$, where $P=(1-q)\cdot\mathcal{N}(0,\sigma^{2})+q\cdot\mathcal{N}(1,\sigma^{2})$ and $Q=(1-q)\cdot\mathcal{N}(0,\sigma^{2})+q\cdot\mathcal{N}(-1,\sigma^{2})$ is a dominating pair of distributions for $\widetilde{\mathcal{M}}$.
Combined with Lemma 2.3 and numerical accountants, this yields tight $(\varepsilon,\delta)$ bounds for DP-SGD with Poisson subsampling under the substitute neighborhood relation.
2.2 Guarantees for the Final Model and Noisy Cyclic Mini-Batch GD
We next consider privacy amplification by iteration (Feldman et al., , 2018), which gives privacy guarantees for the final model of the training iteration. Recent results by Bok et al., (2024) are applicable to the noisy cyclic mini-batch gradient descent (NoisyCGD) for which one epoch of training is described by the iteration:
$\theta_{j+1}=\theta_{j}-\eta\left(\frac{1}{b}\sum\nolimits_{x\in B_{j}}\nabla_{\theta}f(\theta_{j},x)+Z_{j}\right)$
where $Z_{j}\sim\mathcal{N}(0,\sigma^{2}I_{d})$ and the data $D$ is divided into disjoint batches $B_{1},\ldots,B_{k}$, each of size $b$. The analysis by Bok et al., (2024) also considers the substitute neighborhood relation of datasets. A central element for the DP analysis is the gradient sensitivity:
@@ 2.2 Guarantees for the Final Model and Noisy Cyclic Mini-Batch GD
A loss function family $\mathcal{F}$ has a gradient sensitivity $L$ if $\sup_{f,g\in\mathcal{F}}\left\lVert\nabla f-\nabla g\right\rVert\leq L.$
For example, for a family of loss functions of the form $h_{i}+r$, where $h_{i}$’s are $L$-Lipschitz loss functions and $r$ is a regularization function, the sensitivity equals $2L$. We will use the following result to analyse the DP guarantees of NoisyCGD. Recall that a function $f$ is $\beta$-smooth if $\nabla f$ is $\beta$-Lipschitz, and it is $\lambda$-strongly convex if the function $g(x)=f(x)-\frac{\lambda}{2}\left\lVert x\right\rVert_{2}^{2}$ is convex.
@@
Consider $\lambda$-strongly convex, $\beta$-smooth loss functions with gradient sensitivity $L$. Then, for any $\eta\in(0,2/\beta)$, NoisyCGD is $\mu$-GDP for
@@
We could alternatively use the RDP analysis by Ye and Shokri, (2022). However, as also illustrated by the experiments of Bok et al., (2024), the bounds given by Thm. 2.7 lead to slightly lower $(\varepsilon,\delta)$-DP bounds for NoisyCGD. To benefit from the privacy analysis of Thm. 2.7 for NoisyCGD, we add an $L_{2}$-regularization term with a coefficient $\tfrac{\lambda}{2}$. This makes the loss function $\lambda$-strongly convex.
Figure 1: Values of the product of the learning rate $\eta$ and the $L_{2}$-regularization constant $\lambda$ that lead to tighter privacy bounds for the final model using Thm. 2.7, compared to the whole sequence of updates using the DP-SGD analysis. Here $n=6\cdot 10^{4}$, $b=1000$, $\sigma=15.0$ and $\delta=10^{-5}$.
Finding suitable hyperparameter values for the learning rate $\eta$ and regularization parameter $\lambda$ is complicated by two aspects. The larger the regularization parameter $\lambda$ and the learning rate $\eta$ are, the faster the model ‘forgets’ the past updates and the quicker the $\varepsilon$-values converge. This is reflected in the GDP bound of Thm.  2.7 in the constant $c$, which generally equals $\left|1-\eta\lambda\right|$. To benefit from the bounds of Thm.  2.7, the product $\eta\lambda$ should not be too small. Alternatively, when $\eta\lambda$ is too large, the ’forgetting’ affects the model performance. We experimentally observe that the plateauing of the model accuracy and privacy guarantees occurs approximately simultaneously.
Figure 1 illustrates the privacy guarantees of NoisyCGD for a range of values for the product $\eta\lambda$. The $(\varepsilon,\delta)$-DP guarantees given by Thm.  2.7 become smaller than those given by the Poisson subsampled DP-SGD with an equal batch size $b=1000$ when $\sigma=15.0$ and training for 400 epochs. For a given learning rate $\eta$, we can always adjust the value of $\lambda$ to have desirable $(\varepsilon,\delta)$-DP guarantees. To put the values of Fig. 1 into perspective, in experiments we observe that $\eta\lambda=2\cdot 10^{-4}$ is experimentally found to affect the model performance already considerably whereas $\eta\lambda=1\cdot 10^{-4}$ affects only weakly.
Computational advantages of NoisyCGD. NoisyCGD offers significant computational benefits over traditional DP-SGD. The standard DP-SGD requires Poisson sampling where each sample is included in a batch with an independent probability, leading to variable batch sizes. This variability creates practical challenges: For a target physical batch size of 256, one typically needs to set the expected batch size to 230 to handle size variations, resulting in an 11% throughput reduction. Additionally, approximately 4% of batches exceed the physical limit and must be processed separately as stragglers. This further reduces efficiency and increases the engineering complexity of experimental code.
While Chua et al., 2024b proposed truncating oversized batches and accounting for this in the $\delta$ parameter of $(\varepsilon,\delta)$-DP, this approach also impacts throughput. For example, with a physical batch size of 256, $\varepsilon=8$, $\delta=10^{-6}$, and 10 training epochs, the expected batch size must be reduced to 166 according to their theorem. This constitutes a 35% loss of throughput. In contrast, NoisyCGD maintains constant batch sizes (except for possibly the final batch). This enables a simpler implementation in code and higher throughput while preserving privacy guarantees.